OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 6 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now. Ralph's age is _____.

OpenStudy (dan815):

let Age of Ralph be R Let age of Sara be S R=3S currently in 6 years R+6=2(S+6)

OpenStudy (dan815):

3S+6=2S+12 S=6

OpenStudy (dan815):

R=3*6=18

OpenStudy (anonymous):

let age of Sara=x Ralph age=3 x after 6 years age of Sara=x+6 after 6 years age of Ralph=3x +6 by the given condition 3 x+6=2(x+6) find x ,then age

OpenStudy (anonymous):

6 x 2 = 12 12 x 3= 36 so I think the answer is 35 unless I did this wrong

OpenStudy (anonymous):

I mean 36

OpenStudy (anonymous):

Thanks @lexie<3

OpenStudy (anonymous):

no,you are not correct.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 6 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now. Ralph's age is _____. A.6 B.12 C.18 D.24

OpenStudy (anonymous):

Its D

OpenStudy (anonymous):

oh sorry I guess I was wrong

OpenStudy (anonymous):

again it is not correct.

OpenStudy (anonymous):

3x+6=2x+12 3 x-2 x=12-6 x=6 3x=3*6=?

OpenStudy (anonymous):

18?

OpenStudy (anonymous):

yeahhh

OpenStudy (anonymous):

so Ralph's age=?

OpenStudy (anonymous):

18!

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes gimp

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

im freshhh frooom costaaaaaaaa ricaaaaaaaaaaa